# Sequential giant planet formation initiated by disc substructure

### v1

@@ 5 Conclusions
This work demonstrates a scenario of sequential giant planet formation that is triggered by an initial disc substructure. We further extended the model in Lau et al. (2022) by including the effects of planetary gas accretion and gap opening. DustPy is employed to model a protoplanetary disc with micron-sized dust initially and SyMBAp is employed to model the evolution of the planetesimals upon formation.
Consistent with the previous results, planetary cores are formed rapidly from the initial disc substructure, which can then be retained at the migration trap and start gas accretion. The results show multiple (up to three) cores can form and grow into giant planets form in each generation. As the first generation of gas giants has formed and opened a steady gap, the new pressure bump at the outer edge of the planetary gap becomes the next location of planet formation.
In the case of the higher value of $\alpha=5\times 10^{-4}$, the second generation of planet formation occurs about 1 Myr after the first one, where only ice giants are formed instead of gas giants. This case also shows a high planet formation efficiency that more than 85% of the dust beyond 5 au is converted into massive bodies. As the first generation of gas giants effectively prevent dust from flowing through to reach the inner disc, the retained dust is then available for the next generation of planet formation. In the case of the lower value of $\alpha=3\times 10^{-4}$, a larger degree of stochasticity is shown while the general scenario of sequential giant planet formation remains. In both cases, a compact chain of giant planets are formed at the end of the simulations. While the simulations are stopped at 2 Myr, a natural continuation to the model is including the effect of photoevaporation to physically dissipate the disc and stop gas accretion.
Although the formation mechanisms of disc substructure are beyond the scope of this work, further investigations are required to study the possible shape and location produced by physical processes. It is unlikely that any disc substructure can trivially provide the conditions required for planetesimal or planet formation. Also, the parameter space and the number of random simulations in this work are limited by the computational costs. Further code optimization is required to study the statistical effects and to model the diversity of planetary systems. And, planetary gas accretion is still an active field of research. Further investigations specifically on gas accretion are required to model the formation time of the Solar System’s giant planets.

### v2

@@ 5 Conclusions
This work demonstrates a scenario of sequential giant planet formation that is triggered by an initial disc substructure. We further extended the model in Lau et al. (2022) by including the effects of planetary gas accretion and gap opening. We employed DustPy to model a protoplanetary disc initially with micron-sized dust, and SyMBAp was employed to model the evolution of the planetesimals upon formation.
Consistent with the previous results, planetary cores are formed rapidly from the initial disc substructure, which can then be retained at the migration trap and start gas accretion. The results show multiple (up to three) cores can form and grow into giant planets in each generation. As the first generation of gas giants has formed and opened a steady gap, the new pressure bump at the outer edge of the planetary gap becomes the next location of planet formation.
In the case of the higher value of $\alpha=5\times 10^{-4}$, the second generation of planet formation occurs about 1 Myr after the first one, and only ice giants were formed instead of gas giants. This case also shows a high planet formation efficiency where more than 85% of the dust beyond 5 au is converted into massive bodies. As the first generation of gas giants effectively prevent dust from flowing through to reach the inner disc, the retained dust is then available for the next generation of planet formation. In the case of a lower value of $\alpha=3\times 10^{-4}$, a larger degree of stochasticity was shown, while the general scenario of sequential giant planet formation remains. In both cases, a compact chain of giant planets are formed at the end of the simulations. While the simulations were stopped at 2 Myr, a natural continuation to the model would be to include the effect of photoevaporation to physically dissipate the disc and stop gas accretion.
Although the formation mechanisms of disc substructure are beyond the scope of this work, further investigations are required to study the possible shape and location produced by physical processes. It is unlikely that any disc substructure can trivially provide the conditions required for planetesimal or planet formation. Also, the parameter space and the number of random simulations in this work are limited by the computational costs. Further code optimisation is required to study the statistical effects and to model the diversity of planetary systems. And, planetary gas accretion is still an active field of research. Further investigations specifically on gas accretion are required to model the formation time of the Solar System’s giant planets.